OpenStudy (anonymous):

What is the simplified form of –3z^2(z + 2) – 4(z^2 + 1)? A ) –7x^2 + 1 B ) –3z^3 – 4z^2 – 6z – 4 C ) –3z^3 – 2z^2 – 4 D ) –3z^3 – 10z^2 – 4

OpenStudy (anonymous):

well..-3z^2*z=-3z^3 -3z^2*2=-6z^2

OpenStudy (anonymous):

-4*z^2=-4z^2 -4*1=-4

OpenStudy (anonymous):

alll you have to do now is combine like terms

OpenStudy (anonymous):

It's D ) thank you :)

OpenStudy (anonymous):

np